# v-Hub: A Visual-Centric Humor Understanding Benchmark for Video LLMs

### v1

@@ Abstract
AI models capable of comprehending humor hold real-world promise—for example, enhancing engagement in human-machine interactions. To gauge and diagnose the capacity of multimodal large language models (MLLMs) for humor understanding, we introduce v-Hub, a novel visual-centric video humor understanding benchmark. v-Hub comprises a curated collection of minimally verbal short videos, sourced from classic silent films and online resources, and reflecting real-world scenarios where humor can be appreciated purely through visual cues. Each video clip is paired with rich annotations, including captions, descriptions, and explanations, supporting evaluation tasks like caption matching and humor explanation. To broaden its applicability, we further construct an open-ended video QA task, making it readily integrable into existing video understanding benchmarks. We evaluate a diverse set of MLLMs, from specialized Video-LLMs to versatile OmniLLMs that can process audio, covering both open-source and proprietary domains. The experimental results expose the difficulties MLLMs face in comprehending humor from visual cues alone. For example, all models exhibit a marked performance drop on caption matching when moving from text-based to video-based evaluation (without audio). Our findings also demonstrate that incorporating audio helps with video humor understanding, highlighting the informativeness of sound and the promise of integrating richer modalities for complex video understanding tasks.
: Equal Contribution. $\dagger$: Project Lead. ✉: Corresponding Author. Contact: shi_zpeng@sjtu.edu.cn,lihengli@stu.pku.edu.cn,yannzhao.ed@gmail.com.
(c) Visuals+Audio. As the man flips through the pages, cartoon characters gradually appear, accompanied by a distinct melody. First, the dancer’s rhythm and the suona player’s piercing tune, then the cymbal player’s resonant clash, together creating an evolving effect.
(d) Visuals+Audio+Text. A guy messaged his friend that he was making a birthday cake for them. After it was baked and sliced, the inside mimicked their chat bubble layout. The whole scene was made even merrier by the Happy Birthday melody.
@@ 1 Introduction
Humor enriches our daily lives and appears in many forms, from jokes and cartoons to comedies and viral videos. AI models capable of understanding humor hold promise for engaging with humans empathetically (Hampes, 2001; 2010), but perceiving and comprehending humor can be challenging even to humans due to the heavy reliance on nontrivial reasoning, social and cultural contexts, etc (see Figure 1). This, on the other hand, makes humor understanding a promising testbed to evaluate how well state-of-the-art AI models understand humor. Indeed, there has been a line of research centering around gauging the capability of pre-trained large language models (LLMs) for humor understanding (Hessel et al., 2022; Hyun et al., 2023; Ko et al., 2023), but parallel work on multimodal LLMs is still lacking, though they are more naturally suited for understanding multimodal humor.
@@
In this work, we address this gap by investigating humor understanding with multimodal LLMs (MLLMs), focusing specifically on MMLMs that are capable of processing video. We choose video as the primary medium of humor, since it captures nuanced variations and diverse styles, presenting a unique challenge for MLLMs. For example, perceiving the humor in Figure1(d) requires recognizing visual text and the layout of chat bubbles and understanding their temporal and semantic correspondences with the cut surface of the cake slice. While there have been a few benchmarks containing humorous videos (see Table 1), all of them were designed exclusively for the evaluation of LLMs (Ko et al., 2023; Hyun et al., 2023). [^1] They translated videos into language descriptions and performed verbal humor evaluation with LLMs. Moreover, they are limited in that each humor either is dominated solely by spoken language (Hyun et al., 2023) or can be understood only when both the video and linguistic cues are present (Ko et al., 2023), ignoring the fact that humans can understand humor from visual cues alone, exemplified by Charlie Chaplin’s silent comedies.
@@
To address this limitation, we curate a set of visual-centric humorous videos from two complementary sources: Charlie Chaplin’s silent films and user-generated short funny videos. Silent film humor is conveyed through visual cues, but is thematically and culturally constrained due to the scripted performance. To increase diversity, we incorporate user-generated funny short videos from various occasions and cultural backgrounds. We rigorously filtered the videos to retain only those where the humor is primarily visual. Our final dataset consists of videos where humor is derived predominantly from the visual modality (e.g., 99% of all videos), making it visual-centric and more suitable for diagnosing the visual reasoning ability of MLLMs.
@@
To assess how well MLLMs understand humor in video, we create a visual-centric humor understanding benchmark (v-Hub), which consists of three distinct tasks. (1) First, the Caption Matching task challenges MLLMs to align video captions with the corresponding videos. Apart from testing surface-level matching, the task is carefully designed to require an appreciation of nuanced, extended humor. (2) Second, the Humor Explanation task evaluates whether MLLMs can extract humor elements and provide accurate rationales. (3) Finally, the Open-ended QA task evaluate the MLLMs’ fundamental understanding of videos from humor genre across temporal, descriptive, and causal dimensions, broadening the applicability of v-Hub. Together, these tasks provide a comprehensive framework to benchmark MLLMs in visual-centric humor understanding.
@@
We evaluate representative MLLMs from both open- and closed-source domains. Depending on the input modalities, we consider the following three task settings. (1) The Text-Only setting assumes human-level interpretation of video contents and provides detailed human-written descriptions. (2) The Video-Only setting offers only videos (without audio) to assess the ability of MLLMs to derive humor solely from visual cues. (3) We further propose a novel Video+Audio setting that combines visual and auditory signals to determine whether sound cues—such as background music and sound effects—help MLLMs (aka. OmniLLMs) better understand humor. [^2] In this work, audio primarily refers to environmental sound rather than human speech (see Section 2.2).
@@
We empirically find that MLLMs generally perform better with text-only inputs than with video-only inputs (see Table 2). For example, Qwen2.5-VL-72B drops in accuracy from 0.719 to 0.673 on Caption Matching, and Gemini-2.5-Flash from 0.611 to 0.583, under the video-only setting, indicating their struggles in capturing subtle visual cues for humor understanding. Adding audio yields slight improvements across most OmniLLMs. For instance, MiniCPM-2.6-o improves from 0.364 to 0.404 in accuracy on Caption Matching, confirming the effectiveness of the audio modality, though it still lags behind the text-only setting. Overall, our v-Hub presents a new challenge and contributes to a comprehensive evaluation of MLLMs. It exposes their weakness in visual-centric humor understanding, stresses the need for enhancing their visual reasoning capabilities, and highlight the promise of integrating additional modalities like sound for video understanding.
Table 1: Comparison between v-Hub and prior humor video datasets.
Dataset Type Visual-Centric Tasks
Humor Explanation Caption Matching Open-ended QA
NYCC (Hessel et al., 2022) Cartoon [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MUStARD (Castro et al., 2019) Sitcom [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
WITS (Kumar et al., 2022) Sitcom [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UR-FUNNY (Hasan et al., 2019) TED Talks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SMILE (Hyun et al., 2023) Sitcom, TED Talks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ExFunTube (Ko et al., 2023) Short videos [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
v-Hub (ours) Short videos, Silent films [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2 Curating Visual-Centric Humorous Videos
@@ 2.1 Humor Video Sources
Our goal is to collect humorous videos that are visual-centric and illustrate diverse humor. A straightforward approach is to collect humorous clips from silent comedies that are entirely devoid of speech. Though silent films may contain recorded music, sound effect, and few captions, which may contribute to the expression of humor, the humor primarily arises from the visual modality. A major issue with silent film clips is that they have rather narrow themes and employ limited storytelling techniques. To enhance the diversity of humor in our dataset, we further incorporate user-generated short funny videos from the Internet. Specifically, we selected videos from an X account (@humansnocontext) that frequently shares humorous clips with minimal reliance on speech or text-based context. Thus, our dataset comprises humorous videos from two different domains that complement each other (see Figure 2):
@@
  • User-Generated Funny Videos: We reviewed the X user @humansnocontext’s tweets posted between March 28, 2023 and October 12, 2024 and collected 18080 short funny videos.
@@ 2.2 Preprocessing and Filtering
We preprocess and filter the initially collected videos according to duration, appropriateness, and speech reliance, sequentially. (1) Duration: We retain videos ranging from 5 to 60 seconds long. Short clips under 5 seconds generally fail to convey meaningful humor, while clips exceeding 1 minute often rely on dialogue. For silent films, we segment long scenes to isolate individual humorous moments, ensuring that each segment captures the full humor, without becoming too long for generation tasks. (2) Appropriateness: To ensure that the contents of our videos are appropriate, we adhered to the safety objectives outlined in Thoppilan et al. (2022) and excluded videos that violated the established criteria (see details in Appendix B.1). (3) Speech reliance: We minimize reliance on speech. Since there is little to no speech in Charlie Chaplin’s silent films, we primarily focused on user-generated funny videos and employed both manual and automatic approaches to filter out speech-heavy videos (see details in Appendix B.1).
@@ 2.3 Annotation
We recruited eight annotators based on the following criteria: (1) sufficient English proficiency to understand video content, (2) broad cultural knowledge to interpret humor arising from various contexts, and (3) strong observational skills assessed through a qualification test (see Appendix B.2). To ensure consistency, we provided detailed guidelines for each annotation task and created a reference manual for on-demand use. Each video underwent three rounds of annotation to guarantee correctness and thoroughness. We conducted the following primary annotation tasks (see Figure 3 for an example annotation):
@@
    • 2.
      Creative Captions extend beyond the video’s original humor by adding imaginative or novel elements (see the added visual caption in Figure 1(b)).
    The dual-caption annotation supports a comprehensive assessment of humor in video from both comprehension and generation perspectives (see Section 3).
  • Video Description: Annotators were instructed to describe the events in each video without making inferences, focusing only on observable objects, actions, and expressions. After the first annotator completes the video description, subsequent annotators review and refine these descriptions for correctness and completeness.
@@
  • Humor Explanation: Three annotators sequentially create and refine humor explanations by adding missing details, guaranteeing comprehensive coverage of the labeled humor sources through an iterative refinement process.
@@ 2.4 Data Analysis
After all filtering processes, we were left with 960 videos, including 267 silent humor clips from Charlie Chaplin and 693 user-generated short funny videos from the Internet. The total duration of the videos is 4h, and the average duration is around 15s. All of them rely on the visual modality to express humor. We identify two key modalities that dominate the delivery of humor: visuals and audio. Apart from 600 videos (63%) conveying humor primarily via pure visual cues (denoted by ‘Visual’), 92 videos (9%) contain additional linguistic cues in visual form—such as embedded captions and subtitles (denoted by ‘Visual+Text’)—that extend humor, 214 video humor (22%) is enhanced by additional sound that covers non-speech auditory elements, such as background music, sound effects, and character vocalizations (denoted by ‘Visual+Audio’), and 46 videos (5%) convey humor through visuals, sound, and visual text (denoted by ‘Visual+Audio+Text’). The video distribution over the four groups is illustrate in Figure 8 in Appendix C, and more analysis can be found there.
3 v-Hub: a visual-centric humor understanding benchmark
@@ 3.1 Evaluation tasks
  • Open-ended QA. To further assess the fundamental understanding of video content, we generate a set of open-ended question-answer pairs for each video (see details in Appendix D.1). These questions—automatically generated by GPT-4o (Hurst et al., 2024) and manually verified—encompass temporal, descriptive, and causal aspects (Xiao et al., 2021). [^3] There are 62, 675, and 223 QA pairs for temporal, descriptive, and causal questions, respectively. This extends the benchmark beyond humor-specific reasoning, providing a broader assessment of video reasoning skills.

### v2

# v-Hub: A Benchmark for Video Humor Understanding from Vision and Sound

@@ Abstract
AI models capable of comprehending humor hold real-world promise—for example, enhancing engagement in human-machine interactions. To gauge and diagnose the capacity of multimodal large language models (MLLMs) for humor understanding, we introduce v-Hub, a novel video humor understanding benchmark. v-Hub comprises a curated collection of non-verbal short videos, reflecting real-world scenarios where humor can be appreciated purely through visual cues. We pair each video clip with rich annotations to support a variety of evaluation tasks and analyses, including a novel study of environmental sound that can enhance humor. To broaden its applicability, we construct an open-ended QA task, making v-Hub readily integrable into existing video understanding task suites. We evaluate a diverse set of MLLMs, from specialized Video-LLMs to versatile OmniLLMs that can natively process audio, covering both open-source and proprietary domains. The experimental results expose the difficulties MLLMs face in comprehending humor from visual cues alone. Our findings also demonstrate that incorporating audio helps with video humor understanding, highlighting the promise of integrating richer modalities for complex video understanding tasks.
v-Hub: A Benchmark for Video Humor Understanding
from Vision and Sound
[…]
$\dagger$: Project Lead. 🖂: Corresponding Author. Contact: shi_zpeng@sjtu.edu.cn,yannzhao.ed@gmail.com.
(a) Visuals+Audio. As the man flips through the pages, cartoon characters gradually appear, accompanied by a distinct melody. First, the dancer’s rhythm and the suona player’s piercing tune, then the cymbal player’s resonant clash, together creating an evolving effect.
(b) Visuals+Audio+Text. A guy messaged his friend that he was making a birthday cake for them. After it was baked and sliced, the inside mimicked their chat bubble layout. The whole scene was made merrier by the Happy Birthday melody.
@@
Table 1: Humor understanding benchmark comparison. Modality indicates the modalities used for humor understanding, where T, V, and A denote text, visual, and audio, respectively. Our v-Hub is visual-centric, as it contains humor derived predominantly from visual cues (V) and enhanced by environmental sound (A).
Dataset Type Source Average Duration Modality Tasks
Explanation Matching Open-ended QA
NYCC (Hessel et al., 2022) Cartoon New Yorker Caption Contests T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Oogiri-GO (Zhong et al., 2024) Images Oogiri Game T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MUStARD (Castro et al., 2019) Sitcom TV Shows 19 sec. T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
WITS (Kumar et al., 2022) Sitcom TV Shows 17 sec. T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UR-FUNNY (Hasan et al., 2019) Speech TED Talks 20 sec. T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SMILE (Hyun et al., 2024) Sitcom, Speech TV Shows, TED Talks 28 sec. T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ExFunTube (Ko et al., 2023) Short videos Youtube 16 sec. T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
HumorQA (Xie et al., 2024) Surprising videos Youtube 07 sec. T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
v-Hub (ours) Short videos, Silent comedies X, Youtube, Charlie Chaplin’s silent films 14 sec. T, V, A [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 1 Introduction
Humor enriches our daily lives and appears in many forms, from jokes and cartoons to comedies and viral videos. AI models capable of understanding humor hold promise for engaging with humans empathetically (Hampes, 2001, 2010), but perceiving and comprehending humor can be challenging even to humans due to the heavy reliance on nontrivial reasoning, social and cultural contexts, etc (see Figure 1). This, on the other hand, makes humor understanding a promising testbed to evaluate how well state-of-the-art AI models understand humor. Indeed, there has been a line of research centering around gauging the capability of pre-trained large language models (LLMs) for humor understanding from visuals and text (Hessel et al., 2022; Hyun et al., 2024; Ko et al., 2023; Chen et al., 2024), but parallel studies of multimodal LLMs are limited, though they are naturally suited for understanding multimodal humor.
@@
In this work, we address this gap by investigating humor understanding with multimodal LLMs (MLLMs), focusing specifically on MLLMs that can process video. We choose video as the primary medium of humor, since it captures nuanced variations and diverse styles, presenting a unique challenge. For example, perceiving the humor in Figure1(b) requires recognizing visual text (or background music) and the layout of chat bubbles and understanding their temporal and semantic correspondences with the cut surface of the cake slice.
While there have been a few benchmarks on video humor understanding (see Table 1), most of them are designed exclusively for the evaluation of LLMs (Ko et al., 2023; Hyun et al., 2024), [^1] They translated videos into language descriptions and performed verbal humor evaluation with LLMs. and the curated humor is dominated by linguistic cues, ignoring the fact that humans can understand humor from visual cues alone, exemplified by silent comedies. Though Xie et al. (2024) have conducted a humor understanding evaluation of MLLMs, the evaluation is limited in coverage and temporal complexity. Moreover, as an important component of non-verbal humor, environmental sound has been overlooked in their data curation and evaluation.
@@
To address these limitations, we curate a set of visual-centric humorous videos from two complementary sources: Charlie Chaplin’s silent films and user-generated short funny videos. Silent film humor is conveyed through visual cues, but is thematically and culturally constrained due to the scripted performance. To increase diversity, we incorporate user-generated funny short videos from various occasions and cultural backgrounds. We rigorously filtered the videos to retain only those where the humor has no reliance on speech and has a duration longer than 5 seconds. Our final dataset consists of videos where humor is derived predominantly from the visual modality, making it visual-centric and better suited for evaluating MLLMs, including variants that can natively process audio. [^2] In this work, audio primarily refers to environmental sound rather than human speech (see Section 2.2).
@@
To assess how well MLLMs understand humor in video, we create v-Hub, a video humor understanding benchmark. v-Hub offers two typical evaluation tasks for humor understanding. (1) First, the Caption Matching task challenges MLLMs to align video captions with the corresponding videos. (2) Second, the Humor Explanation task evaluates whether MLLMs can extract humor elements and provide accurate rationales. To broaden the applicability of v-Hub, we further construct (3) an Open-ended QA task that evaluates the MLLMs’ fundamental understanding of videos from the humor genre across temporal, descriptive, and causal dimensions. Together, these tasks provide a comprehensive framework to benchmark MLLMs in visual-centric humor understanding.
@@
We evaluate representative MLLMs from both open- and closed-source domains. Depending on the input modalities, we consider the following three task settings. (1) The Text-Only setting assumes human-level interpretation of video contents and provides detailed human-written descriptions. (2) The Video-Only setting offers only videos (without audio) to assess the ability of MLLMs to derive humor solely from visual cues. (3) We further propose a novel Video+Audio setting that combines visual and auditory signals to determine whether sound cues—such as background music and sound effects—help MLLMs (aka. OmniLLMs) better understand humor.
@@
We empirically find that MLLMs generally perform better with text-only inputs than with video-only inputs. For example, Qwen2.5-VL drops in accuracy from 0.726 to 0.666 on Caption Matching, indicating its struggles in capturing subtle visual cues for humor understanding. Adding audio yields notable improvements across most OmniLLMs. For instance, MiniCPM2.6-o improves from 0.362 to 0.442 in accuracy on Caption Matching, though it still lags behind the text-only setting. Overall, v-Hub presents a new challenge and contributes to a comprehensive evaluation of MLLMs. It exposes their weakness in visual-centric humor understanding, stresses the need for enhancing their visual reasoning capabilities, and highlights the promise of integrating additional modalities like sound for video understanding.
2 Visual-Centric Video Humor Curation
@@ 2.1 Humor Video Sources
Our goal is to collect humorous videos that are visual-centric and illustrate diverse humor. A straightforward approach is to collect humorous clips from silent comedies that are entirely devoid of speech. Though silent films may contain recorded music, sound effects, and few captions, which may contribute to the expression of humor, the humor primarily arises from the visual modality. A major issue with silent film clips is that they have rather narrow themes and employ limited storytelling techniques. To enhance the diversity of humor in our dataset, we further incorporate user-generated short funny videos from the Internet. Specifically, we select videos primarily from an X account (@humansnocontext) that frequently shares humorous clips with minimal reliance on speech or text-based context. [^3] Though most videos are sourced from a single X account, the content is sufficiently diverse (see Figure 4), as the account owner collects humorous videos from across the Internet. We also use a subset of YouTube videos from Xie et al. (2024) to diversify video sources. Thus, our dataset comprises humorous videos from two different domains that complement each other (see Figure 2):
@@
  • User-Generated Funny Videos: We reviewed the X user @humansnocontext’s tweets posted between March 28, 2023 and October 12, 2024 and collected 18080 short funny videos. To ensure diverse video sources, we collected 1769 surprising YouTube videos from Xie et al. (2024).
@@ 2.2 Preprocessing and Filtering
We preprocess and filter the initially collected videos according to duration, appropriateness, and speech reliance, sequentially. (1) Duration: We retain videos ranging from 5 to 60 seconds long. Short clips under 5 seconds generally fail to convey meaningful humor, while clips exceeding 1 minute often rely on dialogue. For silent films, we segment long scenes to isolate individual humorous moments, ensuring that each segment captures the full humor, without becoming too long for generation tasks. (2) Appropriateness: To ensure that the contents of our videos are appropriate, we adhered to the safety objectives outlined in Thoppilan et al. (2022) and excluded videos that violated the established criteria (see details in Appendix A.1). (3) Speech Reliance: We minimize reliance on speech. Since there is little to no speech in Charlie Chaplin’s silent films, we primarily focused on user-generated funny videos and employed both manual and automatic approaches to filter out speech-heavy videos (see details in Appendix A.1).
@@ 2.3 Annotation
We recruited eight annotators based on the following criteria: (1) sufficient English proficiency to understand video content, (2) broad cultural knowledge to interpret humor arising from various contexts, and (3) strong observational skills assessed through a qualification test (see Appendix A.2). To ensure consistency, we provided detailed guidelines for each annotation task and created a reference manual for on-demand use. Each video underwent three rounds of annotation to guarantee correctness and thoroughness. We conducted the following primary annotation tasks (see Figure 3 for an example annotation):
@@
    • Creative Captions extend beyond the video’s original humor by adding imaginative or novel elements (see the visual caption in Figure 5(b)).
    The dual-caption annotation supports a comprehensive assessment of humor in video from both comprehension and generation perspectives.
  • Video Description: Annotators were instructed to describe the events in each video, including all the details necessary for understanding the humor, without making inferences, focusing only on observable objects, actions, and expressions. After the first annotator completes the video description, subsequent annotators review and refine the current descriptions for correctness and completeness.
@@
  • Background Knowledge: Annotators determine whether understanding the humor in a video requires background knowledge, which refers to external contextual information that cannot be directly inferred from the videos but is necessary or helpful for understanding humor.
  • Humor Explanation: Three annotators sequentially create and refine humor explanations by adding missing details, guaranteeing comprehensive coverage of the labeled humor sources through an iterative refinement process. In our annotation manual (see Figure 8 in Appendix) for humor explanation, we require that annotators include all the humorous elements they can find, as thoroughly as possible, and explain why these elements make people feel humorous.
@@ 2.4 Data Analysis
Figure 4: Data statistics: (a) Distribution of video lengths; (b) Distribution of visual-centric groups; (c) Distribution of humor types, where Others contains Parody, Satire, Miscellaneous, etc; (d) Distribution of cultural backgrounds, where Others covers Middle Eastern and North African, etc.
Duration.
Diversity.
To show v-Hub contains a variety of humor, we follow Buijzen and Valkenburg (2004) to categorize videos into five categories by humor type (see Figure 4(c)), and follow House et al. (2004) and Ronen and Shenkar (2013) to categorize videos into six categories by cultural background (see Figure 4(d)). The illustrations demonstrate that v-Hub covers a wide range of humor types and cultural backgrounds, thus it is sufficiently diverse.
Visual-centric.
After all filtering processes, we were left with 1218 videos, including 267 silent humor clips from Charlie Chaplin and 951 user-generated short funny videos from the Internet. The total duration of the videos is 4.7h, and the average duration is around 14s. All of them rely on the visual modality to express humor. We identify two key modalities that dominate the delivery of humor: visuals and audio. Apart from 722 videos (59%) conveying humor primarily via pure visual cues (denoted by ‘Visual’), 137 videos (11%) contain additional linguistic cues in visual form—such as embedded captions and subtitles (denoted by ‘Visual+Text’)—that extend humor, 256 video humor (21%) is enhanced by additional sound that covers non-speech auditory elements, such as background music and sound effects(denoted by ‘Visual+Audio’), and 94 videos (8%) convey humor through visuals, sound, and visual text (denoted by ‘Visual+Audio+Text’). The video distribution over the four groups is illustrated in Figure 4(b).
Consensus Evaluation.
For annotations like humor explanation and video description, the second and third annotators reviewed and modified previous annotations to ensure consistency. We employ Krippendorff’s alpha (Krippendorff, 2011) to assess the annotators’ consensus on the humor evaluation, using ‘Low,’ ‘Medium,’ and ‘High’ to indicate the strength of the consensus. In v-Hub, more than 90% of data demonstrated a ‘High’ consensus, while only 0.3% showed a ‘Low’ consensus.
3 v-Hub: A Visual-Centric Video Humor Understanding Benchmark
@@ 3.1 Evaluation Tasks
  • Open-ended QA. To further assess the fundamental understanding of video content, we generate a set of open-ended question-answer pairs for each video (see details in Appendix B.1). These questions—automatically generated by GPT-4o and manually verified—encompass temporal, descriptive, and causal aspects (Xiao et al., 2021). [^4] There are 81, 742, and 395 QA pairs for temporal, descriptive, and causal questions, respectively. This extends the benchmark beyond humor-specific reasoning, providing a broader assessment of video reasoning skills.